Reports 1-1 of 1 Clear search Modify search
AdV-COM (1/√f noise)
mwas - 8:38 Friday 28 November 2025 (68274) Print this report
SR TX offset is useful again

Figure 1. This morning around 6.30 UTC I have put back the SR TX offset to 0.015 the value we have been using from August to October to increase the BNS range (have a diagonal SR misalignment instead of a purely SR TY one). This has worked as in the past, B1p decrease and the BNS range has increase by 5Mpc, but made the frequency noise coupling worse.

Note that BS TX control should be put back to high bandwidth, to have a more robust interferometer, and reduce the flucutation in the frequency noise coupling.

 

Images attached to this report
Comments to this report:
mpinto - 10:12 Friday 28 November 2025 (68275) Print this report

In the past days BS tx loop in full bandwidth has been engaged by hand. This morning I've restored the automatic engagement in the automation (uncommented line 6188 on ITF_LOCK.py,  ACQUIRE_LN3). I'll monitor in the next locks if the loop will eventually misbehave.

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×